# Dynamic planning in hierarchical active inference

### v2

@@ 3.2 A module for iterative transformations
where $\bm{w}_{e}^{(i)}$ is a noise term and $\bm{T}^{(i)}$ is a linear transformation matrix. Backpropagating the extrinsic prediction error $\bm{\varepsilon}_{e}^{(i)}$ leads to simple belief updates:
$\begin{split}{\frac{\partial\bm{g}_{e}^{(i)}}{\partial\bm{\mu}_{e}^{(i-1)}}}^{T}\bm{\varepsilon}_{e}^{(i)}&=\bm{T}^{(i)T}\cdot\bm{\varepsilon}_{e}^{(i)}\\ {\frac{\partial\bm{g}_{e}^{(i)}}{\partial\bm{\mu}_{i}^{(i)}}}^{T}\bm{\varepsilon}_{e}^{(i)}&=\frac{\partial\bm{T}^{(i)}}{\partial\bm{\mu}_{i}^{(i)}}\odot[\bm{\varepsilon}_{e}^{(i)}\cdot\bm{\mu}_{e}^{(i-1)T}]\end{split}$ (29)
@@
Figure 10: (a) In this task, the agent (a 23-DoF human body) has to avoid a moving obstacle; this is possible by defining a repulsive attractor for each extrinsic level. (b) In this task, the agent (a 28-DoF kinematic tree) has to reach four target locations with the extremities of its branches. See [44] for more details. (c) In this task, the agent (the two blue eyes on the left of the top view) has to infer the depth of the red object while fixating on it. The inferred position (and its trajectory) is represented in orange, while the blue trajectory is the center of fixation of the eyes. The bottom two frames show the projection of the object in the eye planes. See [58] for more details.
These equations express the most likely intrinsic and extrinsic states that may have generated the new reference frame. As shown in Figure 9(e), modules are linked through the extrinsic units $\mathcal{U}_{e}^{(i)}$, while $\mathcal{U}_{i}^{(i)}$ performs an internal operation and does not contribute to the hierarchical connectivity. Applying this architecture to kinematics, we can realize a hierarchical model with a multiple-output system, wherein the intrinsic hidden states $\bm{x}_{i}^{(i,j)}$ of a level encode a pair of joint angle and limb length of a single DoF. Iteratively applying roto-translations to an origin (e.g., body-centered) reference frame $\bm{x}_{e}^{(0)}$ – consisting of a Cartesian position and an absolute orientation – will determine the kinematic configuration of the agent in terms of extrinsic coordinates [44]. In addition to this, and differently from PCNs, we can now easily express how every single joint and limb would evolve. Or – which is the same from an active inference perspective – how the agent intends to move its joints and limbs, affording a highly advanced control as demonstrated by the simulations of Figures 10(a) and 10(b). Besides modeling limb dynamics, the IE module can be also applied to non-affine transformations, e.g., perspective projections. As displayed in Figure 10(c), this can be useful for estimating the depth of an object via parallel predictions (e.g., from the eyes or multiple cameras) [58] – a process that active inference casts in terms of target fixation [67]. The modularity of this architecture allows the agent to define dynamic attractors in the 2D projected planes, in the 3D reference frames of the eyes, or as simple vergence-accommodation angles.
@@ 3.3 The self, the objects, and the others
Describing Figure 7(b), we passed over a critical mechanism introduced at the beginning: the characterization of objects for dynamic goal-directed behavior. Recall that the hidden states encode in parallel not only the self but also environmental entities; however, the agent’s model now describes the generative process hierarchically:
$\bm{g}_{e}^{(i)}(\bm{x}_{i}^{(i)},\bm{x}_{e}^{(i-1)})=\begin{bmatrix}\bm{T}_{1}^{(i)}(\bm{x}_{i,1}^{(i)})\cdot\bm{x}_{e,1}^{(i-1)}&\dots&\bm{T}_{N}^{(i)}(\bm{x}_{i,N}^{(i)})\cdot\bm{x}_{e,N}^{(i-1)}\end{bmatrix}$ (30)
@@
For the self, this has a simple explanation, i.e., it just generates, one after the other, the positions of every segment of the kinematic chain depending on its joint angles. Concerning an object, attaching its visual observation to a second extrinsic hidden state of a specific level would lead the latter to encode its Cartesian position. How then should all the previous levels be interpreted? If the generative model maintains the same hierarchy for both the self and the object, backpropagating the extrinsic prediction errors of the second component will eventually infer a potential agent’s configuration in relation to the object. For instance, if the object is linked to the last (i.e., hand) level, this would represent the hand at the object location, while all the previous levels would represent appropriate intermediate positions and angles generating that final location. In other words, the additional factorizations of hidden states and likelihoods do not encode simple target angles or positions as before, but a whole configuration of the self that the agent thinks to be suitable for interacting with an object. Since each level can express a particular dynamics through its hidden causes, the inference of this potential configuration is steered to match both the object’s affordances and the agent’s intentions (e.g., grasping a cup by the handle or with the whole hand). Such inferred beliefs would be subject only to exteroceptive information coming from the objects, while proprioceptive states would be used only to update the agent’s belief over its current configuration.
@@
Figure 11: Interactions between an agent (a 2-DoF light-blue arm), another agent (a 3-DoF purple arm), and an object (a red ball). Generative model of the first agent, composed of three parallel pathways representing the kinematic structures of both agents. For clarity, lateral connections among the model components of each level are not shown. Both the self and the other (or object) in relation to the self depend on the same body-centered reference frame. The first component is subject to both proprioception (yellow dotted lines) and exteroception (red dotted lines), while the other two components are only inferred via exteroception. In this case, the interaction with the second agent just depends on the observation of its last level, leading, e.g., to a hand-shaking action.
@@
Besides modeling object dynamics, this strategy is also useful in multi-agent contexts. One could maintain a hierarchical generative model regarding the kinematic chain of another agent, which would be inferred by exteroceptive observations about all its positions and joint angles, starting from a different body-centered reference frame. As shown in Figure 11, the goal-directed method used for external objects reflects in this case as well: the agent could represent, by a parallel hierarchical pathway, a second agent in relation to itself, expressing a particular kind of interaction (e.g., the hand of the second agent in terms of its own, resulting in a shaking action). These two cases could be interpreted, from a biological perspective, as simulating the functioning of mirror neurons, firing whenever a subject executes a voluntary goal-directed action or when that action is performed by other subjects [100]. Building an internal model with the kinematic chain of the others – both per se and in relation to the self – could be critical to predict (thus, to understand) their intentions. In this view, neural activity results because the agent makes constant predictions over their kinematic structures depending on its hypotheses and the current context [101, 98].
@@
The relationships between the self, the objects, and other agents under active inference may be better understood from the simulation of Figure 12, showing two agents with incompatible goals that depend on each other. Here, both agents are able to infer parallel representations of different kinematic chains, using an effective decomposition of potential and real configurations. Note how one’s current belief is always in between the future state to realize and the actual configuration; this speaks of one of the fundamental aspects of active inference, i.e., that our beliefs never really reflect the state of the affairs of the world, but are always biased toward preferred states – eventually driving action. In general, bodily states, objects, or other agents can be all manipulated in reference frames appropriate for a specific context; this is in line with the hypothesis that cortical columns use object-centered reference frames to encode external elements and more abstract entities [102]. This approach has also some analogies with Active Predictive Coding [103] and Recursive Neural Programs [104], which addressed the part-whole hierarchy learning problem in computer vision by recursively applying reference frame transformations to parts of a scene.
@@
Figure 12: (a) In this task, two agents with different kinematic chains (respectively of 5 and 3 DoF) have two incompatible goals: the first agent (in red) has to reach the elbow of the second agent (in blue), while the second agent has to reach the hand of the first agent. Note that after an initial approaching phase from both, the second agent gradually retracts trying to touch the hand of the first agent. (b) Beliefs of both agents, compared with their actual configurations (displayed in light red for the first agent, and in light blue for the second). The top and bottom graphs respectively show the beliefs of the first and second agents. Specifically, the belief over the actual configuration of the self (in orange or cyan), the belief over the other in relation to the self (in dark red or dark blue), and the belief over the other (in green or purple). Note the belief succession during goal-directed behavior: the potential configuration pointing toward one’s intention (either the elbow or the hand), followed by the actual belief, and then by the actual arm. Also note a slight delay about the inference of the configuration of the other.
@@ 4 The hybrid unit
All the hierarchical models we presented fall short on simulating real-life applications that involve planning actions ahead. In this chapter, we turn to the problem of how to integrate discrete decision-making into continuous motor control. In doing this, we will revisit the basic unit of the first chapter, finally using the second input – the prior over the hidden causes.
@@ 4.1 Dynamic inference by model reduction
Consider the unit in Figure 5(b): recall that some sort of multi-step behavior was achieved, which however depended on higher-level priors about different modalities. In most cases, we need to switch intentions based on lower-level information, affording a more dynamic and less uncertain behavior. Taking a pick-and-place operation as an example, an IE module would be more confident about the success of the first reaching movement if it could rely not just on a tactile belief but also on its intrinsic and extrinsic hidden states. In other words, hidden causes $\bm{v}$ should manage to effectively use both its prior $\bm{\eta}_{v}$ and the dynamics prediction error $\bm{\varepsilon}_{x}$. The latter assumes two different roles depending on which pathway it flows into: the gradient with respect to the hidden states infers the position that is most likely to have generated the current trajectory; conversely, the gradient with respect to the hidden causes infers the most likely combination of gains $v_{m}$, signaling the current status of the trajectory and resulting in a dynamic modulation of intentions. However, this pathway is somewhat problematic because the hidden causes are generated by Gaussian distributions and do not encode proper probabilities. Thus, the gradient $\partial_{\nu}\bm{f}_{x}$ infers just one over many possible combinations of gains, and it makes sense as “inferring the most likely intention to have generated the current trajectory” only in simple contexts and if appropriate assumptions are made. Thus, to implement a correct intention selection, we assume that the hidden causes are generated from a categorical distribution:
@@
where $\bm{H}_{v}$ is an intention preference like $\bm{\eta}_{v}$. In this way, each discrete element of $\bm{v}$ represents the probability that a specific continuous trajectory will be realized.
However, we have now the problem of how to convert discrete hidden causes to continuous hidden states, and vice versa. This can be done via Bayesian model reduction, a technique used to constrain the complexity of full posterior models into simpler and more restrictive (formally called reduced) distributions [61, 62]. Reduced means that the likelihood of some data is equal to that of the full model and the only difference rests upon the specification of the priors – hence, the posterior of a reduced model $m$ can be expressed in terms of the posterior of the full model:
@@
In our case, model reduction means to explain the infinite values a continuous signal may assume by a discrete set of hypotheses. In active inference, models that combine discrete and continuous signals are called hybrid or mixed [48, 59], and a simplified version is shown in Figure 13(a). We can cast this procedure into the usual message passing, where top-down and bottom-up messages between the two domains respectively perform a Bayesian Model Average (BMA) of reduced priors and a Bayesian Model Comparison (BMC) of reduced sensory evidence. In conventional hybrid models, discrete hidden states generate priors for the continuous hidden causes by weighting the probability of each discrete state with a specific reduced prior, which thus represents one among many alternatives that the agent thinks to be the cause of what it is perceiving [48]. Conversely, the hidden causes posterior is compared with such reduced priors to find which one among them could be the best explanation, taking into account their discrete probabilities before observing sensory evidence.
@@
Figure 13: (a) A conventional hybrid architecture, composed of a discrete model at the top and a continuous model at the bottom. For simplicity, we assume that the continuous prior $\bm{\eta}_{v}$ is directly conditioned on discrete hidden states $\bm{s}$. We will cover discrete models in the following section. Here, top-down and bottom-up messages are computed by Bayesian model reduction of some static priors, without the possibility for dynamic planning. (b) Factor graph of the unit with hybrid control. The hidden causes are now generated from a categorical distribution, such that instead of inferring a combination of continuous intention gains, the model correctly infers the most likely intention associated with the current dynamic trajectory. This is done by computing the free energy $\bm{E}_{m}$ corresponding to each intention. Performing a Bayesian model reduction between discrete hidden causes and continuous hidden states lets the agent update its reduced priors at each time step. (c) A simplified graph displaying only the exchange of top-down (red) and bottom-up (blue and green) messages of the hybrid control.
@@
Averaging and comparing continuous alternatives that are fixed and determined a-priori results in the agent’s inability to correctly operate in a changing environment. For instance, if the agent thinks to find an object in one of two locations, it will always reach either one or the other initial guesses, even if the object has been moved to a third location. How then to use the newly available evidence to update our reduced assumptions? By considering the hidden causes as generated from a categorical distribution – as in Equation 31 – we can compare the posterior over the hidden states with the output of the dynamics functions $\bm{f}_{m}$, which thus act as the agent’s reduced priors [69]. More formally, we define $M$ reduced prior probability distributions and a full prior model:
[…]
@@
[…]
@@
$\begin{split}\bm{\eta}_{x}^{\prime}&=\sum_{m}v_{m}\bm{f}_{m}(\bm{x})=\sum_{m}v_{m}\bm{e}_{i,m}\\ \bm{\varepsilon}_{x}&=\bm{\mu}^{\prime}-\bm{\eta}_{x}^{\prime}\end{split}$ (36)
which have the same form of Equations 20 and 21. In fact, the hidden states still perceive a single dynamics prediction error containing the total contribution of every intention. As concerns the bottom-up messages $\bm{l}$, we first write the free energy of each reduced model in terms of the full model. As before, maximizing each reduced free energy makes it approximate the log evidence:
@@
See [61, 62] for a full derivation of BMC under the Laplace assumption, and [69] for more details about the presented approach. Equation 39 is the discrete analogous of Equation 14, but now the bottom-up message encodes a proper discrete distribution and can be used to infer the most likely intention associated with the current dynamic trajectory.
@@
The factor graph of this model, which we call a hybrid unit, is displayed in Figure 13(b), and its inferential process at each continuous instant is better understood if we analyze separately the three different pathways shown in Figure 13(c): (i) during the forward pass, the unit receives a discrete intention prior $\bm{H}_{v}$, performs a BMA with dynamically generated trajectories $\bm{f}_{m}(\bm{x})$ that manipulate the inferred beliefs of every environmental entity $\bm{x}_{n}$, and imposes a prior $\bm{\eta}_{x}^{\prime}$ over the 1st order; (ii) through the first backward pass, the unit accumulates the most likely intention related to the current trajectory by comparing it to the ones generated by the dynamics functions; (iii) in the second backward pass, the unit propagates the dynamics prediction error back to the 0th order to infer the most likely continuous state associated with the trajectory, eventually generating biased observations. After a period $T$, the unit finally computes the difference between the discrete prior and the accumulated evidence, generates a new combination of intentions, and the process starts over.
@@
Figure 14: In this task, the agent has to infer which one among two objects is following. The arm and the two objects move along a circular path, but each with a different velocity. Sequence of time frames (left), and dynamics of hidden causes (right). The two hidden causes are associated with different reaching intentions. As the hand gradually moves away from the first target and approaches the second target, the dynamic accumulation of evidence leads to an increase in the second hidden cause. See [69] for more details.

### v3

@@ 3.2 A module for iterative transformations
where $\bm{w}_{o,e}^{(i)}$ is a noise term, while $\bm{T}^{(i)}$ here indicates a linear transformation matrix, which can express forward kinematics as in the previous examples or non-affine transformations. Backpropagating the extrinsic prediction error $\bm{\varepsilon}_{o,e}^{(i)}=\bm{\mu}_{x,e}^{(i)}-\bm{g}_{e}(\bm{\mu}_{x,i}^{(i)},\bm{\mu}_{x,e}^{(i-1)})$ leads to simple belief updates:
$\begin{split}\partial_{x_{e}^{(i-1)}}\bm{g}_{e}^{(i)T}\bm{\varepsilon}_{o,e}^{(i)}&=\bm{T}^{(i)T}\bm{\varepsilon}_{o,e}^{(i)}\\ \\ \partial_{x_{i}^{(i)}}\bm{g}_{e}^{(i)T}\bm{\varepsilon}_{o,e}^{(i)}&=\partial_{x_{i}^{(i)}}\bm{T}^{(i)}\odot[\bm{\varepsilon}_{o,e}^{(i)}\bm{\mu}_{x,e}^{(i-1)T}]\end{split}$ (29)
@@
Figure 10: (a) In this task, the agent (a 23-DoF human body) has to avoid a moving obstacle; this is possible by defining a repulsive attractor for each extrinsic level. (b) In this task, the agent (a 28-DoF kinematic tree) has to reach four target locations with the extremities of its branches. See [80] for more details. (c) In this task, the agent (the two blue eyes on the left of the top view) has to infer the depth of the red object while fixating on it. The inferred position (and its trajectory) is represented in orange, while the blue trajectory is the center of fixation of the eyes. The bottom two frames show the object projection in the eye planes. See [83] for more details.
These equations express the most likely intrinsic and extrinsic states that may have generated the new reference frame. As shown in Figure 9(e), modules are linked through the extrinsic units $\mathcal{U}_{e}^{(i)}$, while $\mathcal{U}_{i}^{(i)}$ performs an internal operation and does not contribute to the hierarchical connectivity. For motor control, we can realize a hierarchical multiple-output system, wherein the intrinsic hidden states $\bm{x}_{i}^{(i,j)}$ of a level encode a pair of joint angle and limb length of a single DoF. Iteratively applying roto-translations to an origin (e.g., body-centered) reference frame $\bm{x}_{e}^{(0)}$ – consisting of a Cartesian position and an absolute orientation – will determine the kinematic configuration of the agent in terms of extrinsic coordinates [80]. The generative model for the intrinsic and extrinsic units of a single IE module at level $i$ are displayed in Figure 9(f). Compared to the single-level generative model of Figure 7(c) comprising every joint angle of the agent’s body, here we have a dependency between levels encoding distinct joint angles. In particular, an additional term that links the extrinsic unit with the previous level, i.e., $p(\bm{x}^{(i)}_{e}|\bm{x}^{(i)}_{i},\bm{x}_{e}^{(i-1)})$, and multiple distributions $p(\bm{x}_{e}^{(i+1,l)}|\bm{x}_{i}^{(i+1,l)},\bm{x}_{e}^{(i)})$ generating extrinsic predictions for different modules at the next level.
At this point, we can easily express how every single joint and limb would evolve, affording a highly advanced control as demonstrated by the simulations of Figures 10(a) and 10(b). Besides modeling limb dynamics, the IE modules can also be applied to other linear transformations, e.g., perspective projections. As displayed in Figure 10(c), this can be useful for estimating the depth of an object via parallel predictions (e.g., from the eyes or multiple cameras) [83] – a process that active inference casts in terms of target fixation and hypothesis testing [78]. The modularity of this architecture allows the agent to define dynamic attractors in the 2D projected planes, in the 3D reference frames of the eyes, or as simple vergence-accommodation angles. This approach also has some analogies with Active Predictive Coding [91] and Recursive Neural Programs [92], which addressed the part-whole hierarchy learning problem in computer vision by recursively applying reference frame transformations to parts of a scene.
@@ 3.3 The self, the objects, and the others
Describing Figure 7(b), we passed over a critical mechanism introduced at the beginning: the characterization of object affordances. Recall that the hidden states encoded in parallel not only the self but also other environmental entities; however, the agent’s model can now express the generative process hierarchically. This is described by the following likelihood function:
$\bm{g}_{e}^{(i)}(\bm{x}_{i}^{(i)},\bm{x}_{e}^{(i-1)})=\begin{bmatrix}\bm{T}^{(i)}(\bm{x}_{i,0}^{(i)})^{T}\bm{x}_{e,0}^{(i-1)}&\bm{T}^{(i)}(\bm{x}_{i,1}^{(i)})^{T}\bm{x}_{e,1}^{(i-1)}&\dots&\bm{T}^{(i)}(\bm{x}_{i,N}^{(i)})^{T}\bm{x}_{e,N}^{(i-1)}\end{bmatrix}$ (30)
@@
in which every IE module of the hierarchy has distinct factors for the self and every entity. For the self, this has a simple explanation, i.e., it just generates, one after the other, the positions of every segment of the kinematic chain depending on its joint angles. Concerning an object, we could encode its Cartesian position by attaching its visual observation to a second factor of the hidden states at a specific level. If the generative model has the same hierarchy for both the self and the object, backpropagating the extrinsic prediction errors of this second component will eventually infer a potential agent’s configuration in relation to the object, as before. For instance, if the object is linked to the last (i.e., hand) level, this would represent the hand at the object location, while all the previous levels would represent appropriate intermediate positions and angles generating that final location. In other words, the additional factorizations of hidden states and likelihoods reflect here a (deep) hierarchical configuration of the self that the agent thinks to be suitable for object manipulation. Since each level can express some dynamics through its hidden causes, the inference of this potential configuration is steered to match the object’s affordances and the agent’s intentions. As will be shown in the next chapter, this permits flexible adaptation of the kinematic chain depending on the circumstances, as well as representing the hierarchical structures of objects (e.g., tools). The inferred beliefs would be subject only to exteroceptive information from the objects, while proprioceptive states would be used only to update the agent’s belief of its current configuration.
@@
Figure 11: Interactions between an agent (a 2-DoF light-blue arm), another agent (a 3-DoF purple arm), and an object (a red ball). Generative model of the first agent, composed of three parallel pathways representing the kinematic structures of both agents. For clarity, lateral connections among the model components of each level are not shown. Both the self and the other (or object) in relation to the self depend on the same body-centered reference frame. The first component is subject to both proprioception (yellow dotted lines) and exteroception (red dotted lines), while the other two components are inferred via exteroception only. In this case, the interaction with the second agent just depends on the observation of its last level, leading, e.g., to a hand-shaking action.
@@
Besides modeling object affordances, this strategy is also useful in multi-agent contexts. One could maintain a hierarchical generative model regarding the kinematic chain of another agent, which would be inferred by exteroceptive observations about all its positions and joint angles, starting from a different body-centered reference frame. As shown in Figure 11, the goal-directed method used for external objects reflects in this case as well: the agent could represent, by a parallel hierarchical pathway, a second agent in relation to itself, expressing a particular kind of interaction (e.g., the hand of the second agent in terms of its own, resulting in a shaking action). These two cases could be interpreted, from a biological perspective, as simulating the functioning of mirror neurons, firing whenever a subject executes a voluntary goal-directed action or when that action is performed by other subjects [93]. Building an internal model with the kinematic chain of the others – both per se and in relation to the self – could be critical to predict (thus, to understand) their intentions. In this view, neural activity results because the agent makes constant predictions over their kinematic structures depending on its hypotheses and the current context [94, 81].
@@
The relationships between the self, the objects, and other agents under active inference may be better understood from the simulation of Figure 12, showing two agents with incompatible goals that depend on each other. Here, both agents are able to infer parallel representations of different kinematic chains, using an effective decomposition of potential and real configurations. Note how one’s current belief is always in between the intentional state to be realized and the actual configuration; this speaks of one of the fundamental aspects of active inference, i.e., that our beliefs never really reflect the state of the affairs of the world, but are always biased toward preferred states – eventually driving action. In general, bodily states, objects, or other agents can all be manipulated in reference frames appropriate for a specific context; this is in line with the hypothesis that cortical columns use object-centered reference frames to encode external elements and more abstract entities [95].
@@
Figure 12: (a) In this task, two agents with different kinematic chains (respectively of 5 and 3 DoF) have two incompatible goals: the first agent (in red) has to reach the elbow of the second agent (in blue), while the second agent has to reach the hand of the first agent. Note that after an initial approaching phase from both, the second agent gradually retracts trying to touch the hand of the first agent. (b) Beliefs of both agents, compared with their actual configurations (displayed in light red for the first agent, and in light blue for the second). The top and bottom graphs respectively show the beliefs of the first and second agents. Specifically, the belief of the actual configuration of the self (in orange or cyan), the belief of the other in relation to the self (in dark red or dark blue), and the belief of the other (in green or purple). Note the belief succession during goal-directed behavior: the potential configuration pointing toward one’s intentional state (either the elbow or the hand), followed by the current belief, and then by the real arm. Also note a slight delay about the inference of the configuration of the other agent.
@@ 4 The hybrid unit
The continuous-time hierarchical models presented so far lack effective usability in the real world: although they can represent any future trajectory - which implies a degree of planning - they do not possess an explicit model of future states nor can they choose among alternative trajectories. In this chapter, we turn to the problem of how to integrate discrete decision-making into continuous motor control. In doing this, we revisit the basic unit of the first chapter, finally using the second input – the prior over the hidden causes.
Active inference in discrete state-space [96, 97] – generally attributed to the cerebral cortex, especially prefrontal areas [98], along with corticostriatal loops – exploits the structure of Partially Observable Markov Decision Processes (POMDPs) to plan abstract actions over expected sensations. This (active) inference relies on the minimization of the expected free energy, i.e., the free energy that the agent expects to perceive in the future. The expected free energy can be unpacked into two terms resembling the two classical aspects of control theory, exploration and exploitation – which here naturally arise; these respectively correspond to an uncertainty-reducing term, and a goal-seeking term that finds a sequence of actions toward the agent’s prior belief.
Further, so-called mixed or hybrid models [49, 67] combine the potentialities of a discrete model with the inference of continuous signals, allowing robust decision-making in changing environments. While the theory of Bayesian model reduction [99, 100, 101, 102] provides efficient communication between the two models, this unified approach has not enjoyed many practical implementations for the time being [35, 49, 67, 103, 104, 105, 78]. An open issue regards how to deal with highly dynamic environments: hybrid models usually perform a comparison between static priors, limiting the agent to realize, e.g., multi-step reaching movements through fixed positions. In [65], a hybrid model in which the agent’s hypotheses were generated at each time step from the system dynamics allowed to relate continuous trajectories with discrete plans. Besides these more conventional solutions, many other hybrid methods have been proposed recently. One study addressed the problem of realistic robot navigation in active inference, making use of bio-inspired SLAM methods [106]. Other studies proposed how to integrate active inference with imitation learning for autonomous vehicles, using a Dynamic Bayesian Network (DBN) to explain the interactions between an expert agent and dynamic objects [107, 108]. An enhanced DBN consisting of two continuous levels and two discrete levels was used to model the behavior of a UAV at different timescales for assisting wireless communications [109, 110, 111, 112]. A study in robotics combined active inference with behavior trees for reactive action planning in dynamic environments [113]. Finally, a hybrid model based on recurrent switching linear dynamical systems allowed to discover non-grid discretizations of a continuous Mountain Car task. [114].
@@ 4.1 Dynamic inference by model reduction
Recall that in the unit of Figure 5(a), some sort of multi-step behavior was achieved, based on higher-level priors over tactile sensations. In most cases, we need to switch intentions based on lower-level information, affording a more dynamic behavior. Considering a pick-and-place operation, an IE module would be more confident about the success of the first reaching movement if it could rely not just on a tactile belief but also on its kinematic configuration. In other words, hidden causes $\bm{v}$ should manage to effectively use both its prior $\bm{\eta}_{v}$ and the dynamics prediction error $\bm{\varepsilon}_{x}$. The latter assumes two different roles depending on which pathway it flows into: the gradient with respect to the hidden states infers the position that is most likely to have generated the current trajectory, needed for movement; conversely, the gradient with respect to the hidden causes infers the most likely combination of gains $v_{m}$, signaling the current status of the trajectory and resulting in a dynamic modulation of intentions. However, since the hidden causes are generated by Gaussian distributions, they do not encode proper probabilities, and the gradient $\partial_{v}\bm{f}_{x}$ infers just one over many possible combinations of gains. Instead, what we need is to infer the most likely intention to have generated the current trajectory. Thus, to implement a correct intention selection, we assume that the hidden causes are generated from a categorical distribution:
@@
where $\bm{H}_{v}$ is a prior preference like $\bm{\eta}_{v}$. In this way, each discrete element of $\bm{v}$ represents the probability that a specific continuous trajectory will be realized.
Conversion between discrete hidden causes to continuous hidden states (and vice versa) is done via Bayesian model reduction, a technique used to constrain the complexity of full posterior models into simpler and more restrictive (formally called reduced) distributions [100, 101]. Reduced means that the likelihood of some data is equal to that of the full model and the only difference rests upon the specification of the priors; hence, the posterior of a reduced model $m$ can be expressed in terms of the posterior of the full model:
@@
In our case, model reduction means to explain the infinite values of a continuous signal with a discrete set of hypotheses. A simplified version of a hybrid active inference model is shown in Figure 13(a). We can cast this procedure into the usual message passing, where top-down and bottom-up messages between the two domains perform a Bayesian Model Average (BMA) of reduced priors and a Bayesian Model Comparison (BMC) of reduced sensory evidence, respectively. In conventional hybrid models, discrete hidden states generate priors for the continuous hidden causes by weighting a specific reduced prior with the probability of each discrete state; hence, the reduced priors represent alternative hypotheses over the true causes of the sensorium [49]. The posterior over the hidden causes is then compared with such reduced priors to find which one among them could be the best explanation of the environment, taking into account their discrete probabilities before observing sensory evidence.
@@
Figure 13: (a) A conventional hybrid model, composed of a discrete model at the top and a continuous model at the bottom. For simplicity, we assume that the continuous prior $\bm{\eta}_{v}$ is directly conditioned on discrete hidden states $\bm{s}$. We will cover discrete models in the following section. Here, top-down and bottom-up messages are computed by Bayesian model reduction of some static priors, without the possibility for dynamic planning. (b) A simplified graph displaying the exchange of top-down (red) and bottom-up (blue and green) messages of the hybrid control in (c). (c) Factor graph of the unit with hybrid control. The hidden causes are now generated from a categorical distribution, such that instead of inferring a combination of continuous attractor gains, the model correctly infers the most likely intention associated with the current trajectory. This is done by computing the free energy $E_{m}$ corresponding to each intention. Through Bayesian model reduction between discrete hidden causes and continuous hidden states, the agent updates its reduced priors at each time step. Generative model of (c).
@@
Because the agent compares continuous alternatives that are fixed and determined a-priori, it cannot correctly operate in a changing environment. For instance, if the agent thinks to find an object in one of two locations, it will always reach either one or the other initial guesses, even if the object has been moved to a third location. How then to use the newly available evidence to update our reduced assumptions? By considering the hidden causes as generated from a categorical distribution – as in Equation 31 – we can compare the posterior over the hidden states with the output of the dynamics functions $\bm{f}_{m}$, which thus act as the agent’s reduced priors [65]. More formally, we define $M$ reduced prior probability distributions and a full prior model:
$\begin{split}p(\bm{x}^{\prime}|\bm{x},m)&=\mathcal{N}(\bm{f}_{m}(\bm{x}),\bm{\Pi}_{x,m})\\ p(\bm{x}^{\prime}|\bm{x},\bm{v})&=\mathcal{N}(\bm{\eta}_{x}^{\prime},\bm{\Pi}_{x})\end{split}$ (33)
@@
$\begin{split}q(\bm{x}^{\prime}|m)&=\mathcal{N}(\bm{\mu}_{x,m}^{\prime},\bm{P}_{x,m})\\ q(\bm{x}^{\prime})&=\mathcal{N}(\bm{\mu}_{x}^{\prime},\bm{P}_{x})\end{split}$ (35)
@@
$\begin{split}\bm{\eta}_{x}^{\prime}&=\sum_{m}^{M}v_{m}\bm{f}_{m}(\bm{x})=\sum_{m}^{M}v_{m}\bm{e}_{i,m}\\ \bm{\varepsilon}_{x}&=\bm{\mu}_{x}^{\prime}-\bm{\eta}_{x}^{\prime}\end{split}$ (36)
which have the same form of Equations 15 and 17. In fact, the hidden states still perceive a single dynamics prediction error containing the total contribution of every intention. Concerning the bottom-up messages $\bm{l}$, we first write the free energy of each reduced model in terms of the full model. As before, maximizing each reduced free energy makes it approximate the log evidence:
@@
where $\bm{l}=[l_{1},\dots,l_{M}]$. See [100, 101] for a full derivation of BMC under the Laplace assumption, and [65] for more details about the presented approach. Equation 39 is the discrete analogous of Equation 9, but now the bottom-up message encodes a proper discrete distribution and can be used to infer the most likely intention associated with the current trajectory.
@@
The factor graph of this model, which we call a hybrid unit, is displayed in Figure 13(c). Its inference process at each continuous step is better understood if we analyze separately the three different pathways shown in Figure 13(b): (i) during the forward pass, the unit receives a discrete intention prior $\bm{H}_{v}$, performs a BMA with potential trajectories $\bm{f}_{m}(\bm{x})$, and imposes a prior $\bm{\eta}_{x}^{\prime}$ over the 1st order; (ii) through the first backward pass, the unit accumulates the most likely intention related to the current trajectory by comparing it to the ones generated by the dynamics functions; (iii) in the second backward pass, the unit propagates the dynamics prediction error back to the 0th order to infer the most likely continuous state associated with the trajectory, eventually generating biased observations. After a period $T$, the unit finally computes the difference between the discrete prior and the accumulated evidence, generates a new combination of intentions, and the process starts over.
@@
Figure 14: (a) In this perception task, the agent has to infer which one among two objects (a red ball or a grey square) is following. The arm moves along a circular trajectory, while the two objects move linearly in opposite directions. (a) Sequence of time frames (left). The dark blue arrow represents the real agent’s trajectory, while the red and green arrows represent potential trajectories associated with different reaching movements. (b) Dynamics of hidden causes. The two hidden causes $v_{t1}$ and $v_{t2}$ are associated with the two potential trajectories. As the hand moves away from the first target and approaches the second target, the dynamic accumulation of evidence leads to an increase in the second hidden cause. Here, we used a uniform prior $\bm{H}_{v}$, so the hidden causes only express the accumulated log evidence $\bm{l}$. Also, the continuous time $T$ for evidence accumulation was set to $30$ time steps. See [65] for more details.
Now how this mechanism is self-reinforcing: a decision generates some movement, which in turn infers the decision itself to compute the next movement. Inferring a target from body movements is a fundamental aspect often neglected in decision-making studies, which can result in completely different behaviors. In particular, this component produces a commitment effect to the decision taken and stabilizes the actions, avoiding changes of mind that may lead to the loss of valid opportunities in dynamic environments [115, 116]. Further, this kind of dynamic inference has several utilities, e.g., it can be used to infer which one among multiple objects an agent is following – as exemplified in Figure 14 – by generating trajectories for different objects and comparing them with the one perceived [65].
An alternative that produces similar results is the model used in [117] in the context of social exchange. In this hybrid solution, a "student" bird maintains several continuous generative models for every teacher conspecifics. Discrete switching variables placed at the center of these hypotheses infer which teacher bird has generated the perceived birdsong. Learning of the generative models relies on two complementary methods, i.e., Bayesian model average of all possible teacher birds, or Bayesian model selection of a specific bird generating a song.